OpenStudy (anonymous):

Please help....solve using substitution or elimination by addition. 2x + 4y = -8 x + 2y = 4 Thank you will give medals for help

OpenStudy (amorfide):

I would go for substitution as it seems the easiest method due to the coefficients of x and y. re arrange the second equation to make x the subject y=(4-x)/2 substitute y into the first equation so 2x+4y=-8 becomes 2x+4(4-x)/2 =-8 then solve for x to get your y value, substitute your x value into one of the equations.

OpenStudy (anonymous):

Thank you